Two trains are traveling at 18kmph and are 60 km apart. There is fly in the train. it flies at 80kmph. It flies and hits the second train and then it starts to oscillate between the two trains. At one instance when the two trains collide it dies. Distance traveled by the fly when both trains collide is
Found 2 solutions by mananth, sudhanshu_kmr:
Answer by mananth(16949)

distance between two trains = 60 km
speed = 18km/h
travelling towards each other
so speed = 36 km/h
...
time to collide = 60/36= 5/3 hours.
speed of fly = 80 km/h
time it shuttles = 5/3 hours
distance = speed * time = 80*5/3
133.33 kms

Answer by sudhanshu_kmr(1152)


relative velocity of two trains = 36 kmph...
time taken for collision = 60/36 = 5/3 hour

distance traveled by the fly = 5/3 * 80 = 400/3 = 133.333 km
distance traveled by fly = 133.333 km...